# Friends not Foes: Strong Correlation between Inner Super-Earths and Outer Gas Giants

### v1

@@ 6 Effect of Gas Giant Properties on the Correlation
First we consider the impact of gas giant orbital distance on the SE/GG connection. We divide our systems hosting gas giants into two samples, ones with gas giants at 0.13 AU, and another at 310 AU. After doing the same for the W20 sample, we calculate the frequencies of closer and further gas giants as a function of metallicity. Note that for both our sample and W20 we treat multi-gas giant systems the same, counting systems twice if they have gas giants in both bins. In Figure 3 and Table 1, we see that for systems with gas giants 0.13 AU, the positive correlation between super-Earths and gas giants P(GGSE, [Fe/H]0) P(GG[Fe/H]0) gets weaker, decreasing to a significance of 1.7. In contrast, there is a stronger positive correlation between super-Earths and more distant gas giants 310 AU, with P(GGSE, [Fe/H]0) P(GG[Fe/H]0) by 3.0. The increase in SE/GG correlation towards farther giants is in agreement with the result of Rosenthal et al. (2022). What we find additionally is that while such an increase is pronounced around metal-rich hosts, we see no significant SE/GG correlation around metal-poor hosts regardless of gas giant semi-major axis.
How might evidence of a dynamically hot or cold history imprint on the SE/GG connection? We divide our sample into systems that have high eccentricity gas giants ( 0.2), and systems with gas giants that have low eccentricities ( 0.2). For the eccentric gas giant case around metal-rich stars, we see a comparable strength of the positive correlation, P(GGSE, [Fe/H]0) = 19.2 (+4.3 -3.9) P(GG[Fe/H]0) = 7.1 (+2.7 -2.2) at the 2.6 level. However, for dynamically cold metal-rich systems this positive correlation largely disappears, dropping to a significance of 0.9 (Table 1). True to form, metal-poor systems do not exhibit any significant correlation for either dynamically hot or cold systems.
We next explore trends with gas giant multiplicity. Our sample has an average gas giant multiplicity of 1.3, while the W20 sample has an average multiplicity of 1.2. We break gas giant companions into single and multiple gas giant systems, where all gas giants considered fall in the default gas giant parameter space (110 AU, 0.520 M). Table 1 and Figure 3 show that for metal rich systems, selecting for gas giant multiplicity does not produce a difference in the strength of the correlation. While systems with either single or multiple gas giants show a positive correlation, the strength of this correlation is comparable (1.7 vs 1.8). Metal poor systems finally do something different – there is a tentative anti-correlation between super-Earths and multiple gas giants (at the -1.2 level).
@@
We take the calculated P(GGSE) and P(GG) for metal-rich and metal-poor stars as shown in Table 1. For P(SE) we take P(SE) = 303 from Zhu et al. (2018). Around metal-rich stars, P(SEGG, [Fe/H]0) = 56.6 (+22.6 -15.6), and for their metal-poor counterparts, P(SEGG, [Fe/H]0) = 13.0 (+18.9 -5.1). While more than half of metal rich stars hosting cold Jupiters should also host super-Earths, metal poor stars with gas giants should be less likely to host an inner super-Earth than other stars (P(SE)P(SEGG, [Fe/H]0)).
@@ 7 Discussion and Conclusion
Using the largest sample size of systems to-date with both inner super-Earths and outer gas giants, we verify the positive correlation between the two populations that were previously reported by Zhu & Wu (2018b) and Bryan et al. (2019). In particular, our results demonstrate that the correlation is particularly strong for metal-rich ([Fe/H] 0) systems with no statistically significant correlation around metal-poor ([Fe/H] 0) stars, verifying the argument of Zhu (2023) and highlighting the importance of system metallicity in shaping planetary architecture.
@@
In Figure 4, we illustrate the decisive role system metallicity—which we trace with the total initial disk dust mass—plays in establishing inner-outer planet correlation. We reproduce the required dust mass to coagulate a planetary core of pebble isolation mass at each orbital distance as outlined in Chachan & Lee (2023) for 1 host star, turbulent and disk gas accretion rate . We further adopt grain fragmentation velocity of 50 m/s following Kimura et al. (2020, their Figure 13). Comparing with the critical core mass to trigger runaway gas accretion computed in Savignac & Lee (2023) over 1–3.5 AU (under dusty gas accretion), we find that the outer planetary cores will likely nucleate gas giants beyond 3 AU. Furthermore, and most importantly, with higher metallicity (i.e., larger solid reservoir), the spatial extent over which the disk can generate planetary cores enlarges, especially towards wider orbital separations where larger isolation mass and lower planet formation efficiency enforce larger required dust mass. This larger zone of core coagulation and the creation of giant-nucleating massive cores beyond 3 AU around higher metallicity systems therefore explains concurrently the positive correlation between inner super-Earths and outer giants around stars of [Fe/H] 0 and the enhancement of this correlation for giants that are farther away.
@@
In the present paper, we do not find any difference in the strength of inner-outer correlation for systems with single or multiple outer giants. We can see in Figure 4 that at 3 AU, to create 10 core, the total dust mass in the disk needs to be 100 and that the inner cores can be created as long as 20–30 worth of dust can be filtered through. Given the average outer giant multiplicity 1.2–1.3, the number of outer giants is likely at most 2 and so there will be more than enough dust that will drift to the inner region, consistent with the similar level of inner-outer correlation with respect to outer giant multiplicity. In low metallicity systems however, the total mass budget would be quickly depleted with the formation of multiple cores which is in line with the observed anti-correlation around [Fe/H] 0 stars.
@@
An intriguing result from our analysis is the strong inner-outer correlation for gas giants with high eccentricity () and no statistically significant correlation for outer giants with low eccentricity (). While the observed affinity for super-Earth systems to be neighbored by eccentric outer giant(s) could be a reflection of the tendency for metal-rich systems to harbor eccentric giants (see, e.g., Dawson & Murray-Clay, 2013, for Jupiters inside 1 AU), we do not see the latter tendency in our sample, nor in W20 sample, which suggests that the metallicity-eccentricity relation is confined to warm and hot Jupiters. It may be that the more eccentric outer Jupiter aids the formation of inner super-Earths by facilitating core coagulation via orbital instabilities and merger events. Such a hypothesis will be a subject of Paper II where we will explore how the properties of the inner planets (e.g., radius, mass, multiplicity, eccentricities/inclinations, and the innermost orbital period) vary in systems with and without outer giants.
@@
M.L.B. acknowledges support by NSERC, the Heising-Simons Foundation, and by the Connaught New Researcher Award from the University of Toronto. E.J.L. gratefully acknowledges support by NSERC, by FRQNT, by the Trottier Space Institute, and by the William Dawson Scholarship from McGill University.
@@ Appendix A Full sample list and average sensitivities
[…]
@@
GJ 160.2 | 0.690.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | -0.260.1033In the cases where uncertainties on stellar metallicities are not given, we assume an error of 0.1 dex. | 1 | RV | 107 | 8.3 | No | 65 |
@@
HD 102365 | 0.850.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | -0.310.01 | 1 | RV | 149 | 12.5 | No | 72 |
@@
HD 134060 | 1.090.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | 0.140.01 | 2 | RV | 339 | 13.2 | No | 53 |
@@
HD 181433 | 0.780.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | 0.330.13 | 3 | RV | 200 | 13.8 | Yes | 56 |
@@
HD 20003 | 0.880.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | 0.040.02 | 2 | RV | 203 | 13.7 | No | 53 |
@@
HD 20781 | 0.700.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | -0.110.02 | 4 | RV | 237 | 13.7 | No | 53 |
@@
HD 21693 | 0.800.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | 00.02 | 2 | RV | 213 | 11.3 | No | 53 |
@@
HD 31527 | 0.960.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | -0.170.01 | 3 | RV | 268 | 13.9 | No | 53 |
@@
HD 51608 | 0.800.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M | -0.070.01 | 2 | RV | 227 | 13.8 | No | 53 |
@@
Target | GG Mass (M) | GG Semi-major axis (AU) | M (M) | [Fe/H] |

### v2

@@ 6 Effect of Gas Giant Properties on the Correlation
First we consider the impact of gas giant orbital distance on the SE/GG connection. We divide our systems into two samples, those with ‘close’ gas giants, and those with ‘distant’ gas giants. We experiment with four definitions of ‘close’ and ‘distant’: 1) 0.1-3AU and 3-10AU; 2) 0.1-2AU and 2-10AU; 3) 0.3-2AU and 2-10AU; and 4) 0.3-3AU and 3-10AU. This fourth definition matches that of Rosenthal et al. (2022), which explored the impact of gas giant location on the SE/GG correlation with a significantly smaller sample size of super-Earth systems. We divide the R21 sample in the same way. For each of these four distance combinations, we calculate the frequencies of closer and further gas giants as a function of metallicity. Note that for both our sample and R21 we treat multi-gas giant systems the same, counting systems twice if they have gas giants in both bins. For all four sets of semi-major axis ranges for ‘close’ and ‘distant’ gas giants, we find that the positive correlation between super-Earths and close gas giants around metal-rich stars P(GGSE, [Fe/H]0) P(GG[Fe/H]0) is slightly weaker than the positive correlation between super-Earths and distant gas giants by 0.2-0.4, but still significantly positive. This runs counter to the tentative findings from Rosenthal et al. (2022) that suggest close gas giants suppress super-Earth formation in their small sample of 28 super-Earth systems. We additionally find that we see no significant SE/GG correlation around metal-poor hosts regardless of gas giant semi-major axis.
How might evidence of a dynamically hot or cold history imprint on the SE/GG connection? We divide our sample into systems that have high eccentricity gas giants ( 0.2), and systems with gas giants that have low eccentricities ( 0.2). For the eccentric gas giant case around metal-rich stars, we see a slightly stronger strength of the positive correlation, P(GGSE, [Fe/H]0) = 20.1 (+4.5 -4.0) P(GG[Fe/H]0) = 7.1 (+1.5 -1.3) at the 3.0 level. However, for dynamically cold metal-rich systems this positive correlation substantially diminishes, dropping to a significance of 1.3 (Table 1). True to form, metal-poor systems do not exhibit any significant correlation for either dynamically hot or cold systems.
We next explore trends with gas giant multiplicity. Our sample has an average gas giant multiplicity of 1.3, while the R21 sample has an average multiplicity of 1.1. We break gas giant companions into single and multiple gas giant systems, where all gas giants considered fall in the default gas giant parameter space (110 AU, 0.520 MJup). Table 1 and Figure 4 show that for metal rich systems, the positive correlation gets slightly stronger when there are multiple gas giants (going from 1.7 for single GG to 2.2 for multi GG), whereas the correlation gets smaller and possibly becomes an anti-correlation for metal-poor systems with multiple gas giants (going from 0.5 for single GG systems to -0.3 for multi-GG systems.
@@
We take the calculated P(GGSE) and P(GG) for metal-rich and metal-poor stars as shown in Table 1. For P(SE) we take P(SE) = 303 from Zhu et al. (2018). Around metal-rich stars, P(SEGG, [Fe/H]0) = 61.2 (+18.4 -12.7), and for their metal-poor counterparts, P(SEGG, [Fe/H]0) = 21.6 (+26.3 -8.7) (Figure 5). While more than half of metal rich stars hosting cold Jupiters should also host super-Earths, metal poor stars with gas giants might be less likely to host an inner super-Earth than other stars (P(SE)P(SEGG, [Fe/H]0)).
@@ 7 Discussion and Conclusion
Using the largest sample size of systems to-date with both inner super-Earths and outer gas giants, we verify the positive correlation between the two populations that were previously reported by Zhu & Wu (2018) and Bryan et al. (2019). In particular, our results demonstrate that the correlation is particularly strong for metal-rich ([Fe/H] 0) systems with no statistically significant correlation around metal-poor ([Fe/H] 0) stars, verifying the argument of Zhu (2023) and highlighting the importance of system metallicity in shaping planetary architecture.
@@
In Figure 6, we illustrate the decisive role system metallicity—which we trace with the total initial disk dust mass—plays in establishing inner-outer planet correlation. We reproduce the required dust mass to coagulate a planetary core of pebble isolation mass at each orbital distance as outlined in Chachan & Lee (2023) for 1 host star, turbulent and disk gas accretion rate . We further adopt grain fragmentation velocity of 50 m/s following Kimura et al. (2020, their Figure 13; the result also holds for lower fragmentation velocities used in the original work of ()). Comparing with the critical core mass to trigger runaway gas accretion computed in Savignac & Lee (2023) over 1–3.5 AU (under dusty gas accretion), we find that the outer planetary cores will likely nucleate gas giants beyond 3 AU. Furthermore, and most importantly, with higher metallicity (i.e., larger solid reservoir), the spatial extent over which the disk can generate planetary cores enlarges, especially towards wider orbital separations where larger isolation mass and lower planet formation efficiency enforce larger required dust mass. This larger zone of core coagulation and the creation of giant-nucleating massive cores beyond 3 AU around higher metallicity systems therefore explains concurrently the positive correlation between inner super-Earths and outer giants around stars of [Fe/H] 0 and the slight enhancement of this correlation for giants that are farther away. The existence of giants at closer-in distances and their similarly positive correlation with the inner planets around metal-rich stars could be explained by a variation in the critical core mass owing to different opacity sources (see, e.g., Piso et al., 2015; Savignac & Lee, 2023).
@@
In the present paper, we find a slight difference in the strength of inner-outer correlation for systems with single vs. multiple outer giants, weakly favoring the latter. We can see in Figure 6 that at 3 AU, to create 10 core, the total dust mass in the disk needs to be 100 and that the inner cores can be created as long as 20–30 worth of dust can be filtered through. Given the average outer giant multiplicity 1.1–1.3, the number of outer giants is likely at most 2 and so there will be more than enough dust that will drift to the inner region, consistent with the similar level of inner-outer correlation with respect to outer giant multiplicity. In low metallicity systems however, the total mass budget would be quickly depleted with the formation of multiple cores which is in line with the observed anti-correlation around [Fe/H] 0 stars.
@@
An intriguing result from our analysis is the strong inner-outer correlation for gas giants with high eccentricity (). While the observed affinity for super-Earth systems to be neighbored by eccentric outer giant(s) could be a reflection of the tendency for metal-rich systems to harbor eccentric giants (see, e.g., Dawson & Murray-Clay, 2013, for Jupiters inside 1 AU), we do not see the latter tendency in our sample, nor in R21 sample, which suggests that the metallicity-eccentricity relation is confined to warm and hot Jupiters. It may be that the more eccentric outer Jupiter aids the formation of inner super-Earths by facilitating core coagulation via orbital instabilities and merger events. Such a hypothesis will be a subject of Paper II where we will explore how the properties of the inner planets (e.g., radius, mass, multiplicity, eccentricities/inclinations, and the innermost orbital period) vary in systems with and without outer giants.
@@
We thank Eugene Chiang, Heather Knutson, Yanqin Wu, Wei Zhu, and the referee for helpful conversations and feedback. M.L.B. acknowledges support by NSERC, the Heising-Simons Foundation, and by the Connaught New Researcher Award from the University of Toronto. E.J.L. gratefully acknowledges support by NSERC, by FRQNT, by the Trottier Space Institute, and by the William Dawson Scholarship from McGill University.
@@ Appendix A Full sample list and average sensitivities
Target | M⋆ (M⊙) | [Fe/H] | Npl | Disc. Method | Nobs | Baseline (yrs) | GG Companion | Ref. |
@@
GJ 160.2 | 0.690.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | -0.260.1033In the cases where uncertainties on stellar metallicities are not given, we assume an error of 0.1 dex. | 1 | RV | 107 | 8.3 | No | 65 |
@@
HD 102365 | 0.850.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | -0.310.01 | 1 | RV | 149 | 12.5 | No | 72 |
@@
HD 134060 | 1.090.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | 0.140.01 | 2 | RV | 339 | 13.2 | No | 53 |
@@
HD 181433 | 0.780.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | 0.330.13 | 3 | RV | 200 | 13.8 | Yes | 56 |
@@
HD 20003 | 0.880.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | 0.040.02 | 2 | RV | 203 | 13.7 | No | 53 |
@@
HD 20781 | 0.700.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | -0.110.02 | 4 | RV | 237 | 13.7 | No | 53 |
@@
HD 21693 | 0.800.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | 00.02 | 2 | RV | 213 | 11.3 | No | 53 |
@@
HD 31527 | 0.960.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | -0.170.01 | 3 | RV | 268 | 13.9 | No | 53 |
@@
HD 51608 | 0.800.1044In the cases where uncertainties on stellar mass are not given, we assume an error of 0.1 M☉ | -0.070.01 | 2 | RV | 227 | 13.8 | No | 53 |
@@
Target | GG Mass (MJup) | GG Semi-major axis (AU) | M⋆ (M⊙) | [Fe/H] |